OpenStudy (anonymous):

solve for the variable. 4.1b = -20.5

OpenStudy (anonymous):

b = -20.5/4.1 = -5 is the answer...

OpenStudy (anonymous):

b=-20.5/4.1 b=-5

OpenStudy (anonymous):

thanks guys (: (: (: